OpenStudy (anonymous):

rationalize the denominator and simplify the answer: sqrt x/ sqrt x - sqrt 7

OpenStudy (anonymous):

ok multiply top and bottom by \[\sqrt{x}+\sqrt{7}\]

OpenStudy (anonymous):

the denominator will be \[(\sqrt{x}-\sqrt{7})(\sqrt{x}+\sqrt{7})=x-7\]

OpenStudy (anonymous):

then numerator will be whatever you get when you multiply out

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

\[\sqrt{x}(\sqrt{x}+\sqrt{7})=x+\sqrt{7x}\]